OpenStudy (anonymous):

Which ratio correctly compares 60 mm to 45cm? A. 4 : 3 B. 2 : 15 C. 40 : 3 D. 2 : 150

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

heres what you do

OpenStudy (anonymous):

put 60/45 in a calculator and turn that into a fraction

OpenStudy (anonymous):

and you will have your answer

OpenStudy (anonymous):

So the answer should be 2 @taylorz115820

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

@taylorz115820 ?

OpenStudy (anonymous):

what do you mean by 2?

OpenStudy (anonymous):

Oh i ment to say a

OpenStudy (anonymous):

yes that is right good job

OpenStudy (anonymous):

Ok thats for the help

OpenStudy (anonymous):

Thanks is that all you need?

OpenStudy (anonymous):

I need help on one or two more problems

OpenStudy (anonymous):

Ok what are they?

OpenStudy (anonymous):

Which value is equal to 7.3% written as a decimal. A. 0.73 B. 0.0073 C. 0.073 D. 73.0

OpenStudy (anonymous):

All you have to do is divide the 5 by 100 so you would do \[\frac{ 7.3 }{ 100 }\]

OpenStudy (anonymous):

So the answer should be A or C

OpenStudy (anonymous):

Not A or C just C

OpenStudy (anonymous):

Ok thanks for the help

OpenStudy (anonymous):

Yw

OpenStudy (anonymous):

no, in the first 2:15

OpenStudy (anonymous):

How reduce 60/45

OpenStudy (anonymous):

thats 4/3

OpenStudy (anonymous):

60mm=60/10=6cm

OpenStudy (anonymous):

6:45or2:15

OpenStudy (anonymous):

its 60 not 6

OpenStudy (anonymous):

oh shiz you right

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

1 cm=10mm

OpenStudy (anonymous):

I didn't even notice that nice catch

OpenStudy (anonymous):

Thank yall for the help